Use Hess’s law to calculate the heat of reaction for the following:    

2 C6H6(g) + 15 O2(g)   🡪 12 CO2(g)    + 6 H2O(g)




1
Expert's answer
2021-06-08T13:36:19-0400

We know the following heat of reactions:


(i) C(s) + 3 H2(g) → C6H6(l) ; ΔH = +45.9kJ 


(ii) H2​(g) + 1/2 O2​(g) → H2O(l) ; ΔH = −285.9kJ
(iii) C(s) + O2​(g) → CO2(g) ; ΔH = −393.5kJ


And we have to find heat of reaction of this one:


(iv) C6H6(l) + 15/2 O2(g) → 6 CO2​(g) + 3 H2O(l)


Reverse equation (i)

(v) C6​H6(l) → 6 C(s) + 3 H2(g) ; ΔH = −45.9kJ

Multiply equation (ii) with 3.

(vi) 3 H2(g) + 3/2 O2(g) → 3 H2O(l) ; ΔH = −285.9 × 3 = −857.7kJ

Multiply equation (iii) with 6.

(vii) 6 C(s) + 6 O2(g) → 6 CO2(g) ; ΔH = −393.5 × 6 = −2361kJ

Add equations (v), (vi) and (vii)

(iv) C6H6(l) + 15/2 O2(g) → CO2(g) + 3 H2O(l) ; ΔH = −45.9 − 857.7 − 2361 = −3264.6kJ


The enthalpy of combustion of benzene is −3264.6kJ.